OpenStudy (anonymous):

Write the equation for the following situation: In one U.S. town, the annual consumption, s, of sugar (in kg per person) can be estimated by adding 30 to half the years, t, since 1975. Then graph the equation and use the graph to determine the sugar consumption in the year 1991.

OpenStudy (anonymous):

Let the year be Y, then s= 30 + (Y-1975)/2. Note that t = Y-1975 This is a straight line with positive slope 1/2 and S-intercept -(1975-60)/2, Y-intercept (1975-60)

OpenStudy (anonymous):

so t-intercept is +60

OpenStudy (anonymous):

consumption in 1991 is 30 + (1991-1975)/2